OpenStudy (anonymous):

The weather in Vancouver in December is fairly constant. Records indicate that the low temperature for each day of the month tend to have a uniform distribution over the interval −5∘ to 3∘C. (Note that a uniform distribution is one whose probability density function is defined to be constant over the given interval.) A business man arrives on a randomly selected day in December. (a) What is the probability that the low temperature will be above 0∘? (b) What is the probability that the low temperature will be between −2∘ and 1∘? (c) What is the expected low temperature? -1

OpenStudy (anonymous):

I only got the part c right.I don't get why part a is not 1/3 and part b 4/9?

OpenStudy (dumbcow):

there are 8 degrees in the interval a) there are 3 degrees above 0 in the interval probability = 3/8 b) there are 3 degrees between -2 and 1 probability = 3/8

OpenStudy (s):

Hmm for a) i thought it's 3/9 = 1/3 as well as you.. b) i think between -2 and 1 i think is NOT including -2 and 1 so we basically need to find the probability of it being -1 or 0.. which is 2 degrees out of 9.. 2/9

OpenStudy (s):

Isn't there supposed to be 9 degrees? -5, -4, -3, -2, -1, 0, 1, 2, 3?

OpenStudy (dumbcow):

yes but the range is 8, so the temperature is only distributed over and interval of 8 degrees even though its 9 integer degrees when you include the endpoints

OpenStudy (anonymous):

oh i see!!Thanks!!

OpenStudy (s):

Hmm ok, makes sense )

OpenStudy (dumbcow):

S, think continuous not discrete temperatures :)

OpenStudy (s):

Yea, i see now =)